OpenStudy (anonymous):

Solve for A: B=5/7 (A-11)

OpenStudy (anonymous):

so 7(A-11)B=5

OpenStudy (anonymous):

so A-11=5/7B

OpenStudy (anonymous):

hence A=5/7B +11

OpenStudy (anonymous):

5/7 is in equation form

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

5 0ver 7 5/7

OpenStudy (anonymous):

and? whats the problem with that

OpenStudy (anonymous):

ooppss.. sorry.. so the answer is 5/7B+11?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks madam :)

OpenStudy (anonymous):

fan me